Let m1, m2 be positive integers with gcd(m_{1}, m_{2}) = d . Prove that the system of congruences x\equiv r_{1}(mod m_{1}), x\equiv r_{2}(mod m_{2}) has a solution if and only if r_{1} \equiv r_{2}(mod d) .
Answer #1

we have:

x \equiv r1 (mod m1) or we can write x = r1 + m1*A where a is an integer

Similarly

x \equiv r2(mod m2) or we can write x= r2 + m1 * B where B is some integer

What we need to prove is

r1 \equiv r2(mod d) or r1 = r2+d*(C) where C is some integer

here

x = m1*A + r1 ...................eq1

x = m2*B + r2...................eq2

Subtracting equation 2 from 1, we get

0 = m1*A -m2*B + r1 - r2

or

r1 = r2 - (m1*A -m2*B) .......................................eq3

Now we also have : gcd(m1 , m2 ) = d

thus two numbers m1 and m2 can be expressed as a integral multiple of d as

m1 = P * d where p is an integer.............................................eq4

and

m2 = Q * d where q is an integer................................................eq5

and also since , d divides m1 and m2 , it also divides m1-m2.

Thus

from eq 3 , eq4 and eq 5, we get

r1 = r2 - (P*d*A - Q*d*B)

or

r1 = r2 - d * (P*A - Q*B)

or

r1 = r2 + Z where Z = -1 * d * (P*A - Q*B)

thus

r1 = r2 + Z where z is d*(P*A - Q*B) .

which is what we have to prove that is

r1 = r2 + d*C where C = (P*A - Q*B).

Hence proved